A 4 kg block moving right at 6 m/s collides elastically with a 2 kg block moving right at 3 m/s. Calculate the final kinetic energy.
1
Expert's answer
2020-11-23T10:28:42-0500

For elastic collisssions:


"K_i=K_f\\K_f=0.5m_1v_1^2+0.5m_2v_2^2\\K_f=0.5((4)(6)^2+(2)(3)^2)=81\ J"